OpenStudy (anonymous):

A ball has a diameter of 8 inches. What is its surface area? (Use 3.14 for π.) 200.96 in.2 803.88 in.2 401.92 in.2 100.48 in.2

jimthompson5910 (jim_thompson5910):

Use the formula SA = 4pi*r^2 where SA = surface area pi = 3.14 (roughly) r = radius

jimthompson5910 (jim_thompson5910):

keep in mind that r = d/2 r = radius d = diameter

OpenStudy (anonymous):

so u do 8X3.14

jimthompson5910 (jim_thompson5910):

8 isn't the radius

jimthompson5910 (jim_thompson5910):

the radius is 8/2 = 4, so r = 4

jimthompson5910 (jim_thompson5910):

SA = 4pi*r^2 SA = 4*3.14*4^2 SA = ??

OpenStudy (anonymous):

100.48

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

4^2 does NOT mean 4 times 2

jimthompson5910 (jim_thompson5910):

4^2 means "4 squared"

OpenStudy (anonymous):

200.96

jimthompson5910 (jim_thompson5910):

better